OpenStudy (anonymous):

HELP 3. The formula to find the surface area of a sphere is to take the radius of the sphere and square it and multiply that amount by pi and then multiply that by 4. This is written algebraically as S=4πr2, where S is the surface area, r is the radius, and the value of π is 3.14. Rewrite the equation to solve for the radius of the sphere if you know the sphere’s surface area. Then estimate the radius of a sphere given a surface area of 500 meters2. HELP

OpenStudy (raden):

4 pi r^2 = 500 4 * 3.14 r^2 = 500 12.56 r^2 = 500 r^2 = 500/12.56 = 39.8 take the squareroot, then r = sqrt(39.8) = ....

OpenStudy (anonymous):

thanks r u sure about this

OpenStudy (raden):

1000 % be sure, lol

OpenStudy (anonymous):

hahahaha thanks

OpenStudy (raden):

welcome :)